A beam of light enters water at an angle of 60 degrees relative to the normal of the water surface . Find the angle of refraction?



Light passes from air into water . If the angle of refraction is 20° ,what is the index of refraction of the plastic?

1
Expert's answer
2022-04-20T13:46:47-0400

1.

"\beta=\arcsin\frac{\sin\alpha}n=40.5°."

2.

"n=\frac 1{\sin\beta}=2.9."